# Model Selection for Off-policy Evaluation: New Algorithms and Experimental Protocol

### v2

@@ B.2 Sign-flip Average Bellman Error
$~{}\mathbb{E}_{\mu}[|Q_{i}-\mathcal{T}_{M^{\star}}^{\pi}Q_{i}|]$ (15)
$=$ $~{}\mathbb{E}_{\mu}[\textrm{sgn}(Q_{i}(s,a)-(\mathcal{T}^{\pi}Q_{i})(s,a))(Q_{i}-\mathcal{T}^{\pi}Q_{i})]$ (16)
$=$ $~{}\mathbb{E}_{\mu}[\textrm{sgn}(Q_{i}-\mathcal{T}^{\pi}Q_{i})(Q_{i}(s,a)-r-\gamma Q_{i}(s^{\prime},\pi))]$ (17)
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[Q_{\widehat{i}}(s,\pi)\right]\leq 4\cdot\mathcal{C}_{\infty}^{\pi}\cdot V_{\max}\sqrt{\frac{\log\left(2m/\delta\right)}{n}}.$ (19)
@@ C.1 Proof of Theorem 1
$\ell(\theta)\vcentcolon=A\theta-b\in\mathbb{R}^{d},$ (20)
^$\widehat{\ell}(\theta)\vcentcolon=\widehat{A}\theta-\widehat{b}\in\mathbb{R}^{d}$ (21)
@@
$\widehat{\theta}\vcentcolon={}\operatorname*{arg\,min}_{\theta\in\Theta}{}\|\widehat{\ell}(\theta)\|_{\infty}.$ (22)
@@
$\|Q^{\pi}(\cdot)-\phi^{\top}(\cdot)\widehat{\theta}\|_{\infty}$ $=\|\phi^{\top}(\cdot)(\widehat{\theta}-\theta^{\star})\|_{\infty}$ (23)
$=\|\phi^{\top}(\cdot)A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}(A\widehat{\theta}-b)\|_{\infty}$ (24)
$=\|\phi^{\top}(\cdot)A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\ell(\widehat{\theta})\|_{\infty}$ (25)
$=\max_{s,a}\lvert\phi^{\top}(s,a)A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\ell(\widehat{\theta})\rvert$ (26)
$\leq\left(\max_{s,a}{}\left\|\phi^{\top}(s,a)\right\|_{2}\right)\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\|\ell(\widehat{\theta})\|_{2}$ (27)
$\leq\sqrt{d}B_{\phi}\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\|\ell(\widehat{\theta})\|_{\infty}$ (28)
Next, we control the $\ell(\widehat{\theta})$ term. In the sequel we will establish via concentration that
@@
$\|\ell(\widehat{\theta})\|_{\infty}$$\leq\|\widehat{\ell}(\widehat{\theta})\|_{\infty}+\varepsilon_{\mathsf{stat}}$ (30)
^$\leq\|\widehat{\ell}(\theta^{\star})\|_{\infty}+\varepsilon_{\mathsf{stat}}$ (31)
$\leq\|\ell(\theta^{\star})\|_{\infty}+2\cdot\varepsilon_{\mathsf{stat}}$ (32)
$=2\cdot\varepsilon_{\mathsf{stat}},$ (33)
@@
$\|Q^{\pi}-\phi^{\top}\widehat{\theta}\|_{\infty}$ $\leq 2\sqrt{d}B_{\phi}\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\varepsilon_{\mathsf{stat}}$ (34)
$=6\sqrt{d}\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\max\{R_{\max},B_{\phi}\}^{2}\cdot\sqrt{\frac{\log(2d\lvert\Theta\rvert\delta^{-1})}{n}},$ (35)
@@
$B^{\pi}(s,a,s^{\prime};{}\theta)\vcentcolon={}\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)$ (36)
@@
$\phi_{j}(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)$ (37)
$\leq$ $~{}\|\phi(s,a)\|_{\infty}\left(\|\phi(s,a)\|_{2}\|\theta\|_{2}+\gamma\|\phi(s^{\prime},\pi)\|_{2}\|\theta\|_{2}+R_{\max}\right)$ (38)
$\leq$ $~{}\max_{s,a}\|\phi(s,a)\|_{2}\left(\max_{s,a}\|\phi(s,a)\|_{2}\|\theta\|_{2}+\gamma\max_{s,a}\|\phi(s,a)\|_{2}\|\theta\|_{2}+R_{\max}\right)$ (39)
$\leq$ $~{}(1+\gamma)B_{\phi}^{2}+R_{\max}B_{\phi}$ (40)
$\leq$ $~{}3\max\{B_{\phi},R_{\max}\}^{2}.$ (41)
@@
$\left\lvert\operatorname{\mathbb{E}}_{\mu}\left[\phi^{j}(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi^{j}(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]\right\rvert\leq 3\max\{B_{\phi},R_{\max}\}^{2}\sqrt{\frac{2\log(d\lvert\Theta\rvert\delta^{-1})}{n}}=\varepsilon_{\mathsf{stat}},$ (42)
@@
$\left\|\ell(\theta)-\widehat{\ell}(\theta)\right\|_{\infty}$ $=\left\|\operatorname{\mathbb{E}}_{\mu}\left[\phi(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)\right]\right\|_{\infty}$ (43)
$=\left\|\operatorname{\mathbb{E}}_{\mu}\left[\phi(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]\right\|_{\infty}$ (44)
$\leq\left\|\mathbf{1}\cdot\varepsilon_{\mathsf{stat}}\right\|_{\infty}$ (45)
$\leq\varepsilon_{\mathsf{stat}}.$ (46)
@@ C.2 Proof of Theorem 2
$\lvert J_{M^{\star}}(\pi)-\mathbb{E}_{s\sim d_{0}}[Q_{\hat{i}}(s,\pi)]\rvert=\lvert\mathbb{E}_{s\sim d_{0}}[Q^{\pi}(s,\pi)]-\mathbb{E}_{s\sim d_{0}}[Q_{\hat{i}}(s,\pi)]\rvert\leq\|Q^{\pi}(\cdot)-Q_{\hat{i}}(\cdot)\|_{\infty}.$
@@
$\|Q^{\pi}(\cdot)-Q_{\hat{i}}(\cdot)\|_{\infty}$ $=\|\phi^{\top}_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{\infty}$ (47)
$=\|\phi^{\top}_{\hat{i},i^{\star}}A^{-1}_{\hat{i},i^{\star}}A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{\infty}$ (48)
$=\max_{s,a}\lvert\phi^{\top}_{\hat{i},i^{\star}}(s,a)A^{-1}_{\hat{i},i^{\star}}A_{\hat{i},i^{\star}}(e_{2}-e_{1})\rvert$ (49)
$\leq\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{2}$ (50)
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{\infty}$ (51)
$=\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}e_{1}-b_{\hat{i},i^{\star}}\|_{\infty}$ (52)
$=\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|\ell_{\hat{i},i^{\star}}\|_{\infty}$ (53)
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\|\hat{\ell}_{\hat{i},i^{\star}}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$ (54)
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\hat{\ell}_{\hat{i},j}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$ (55)
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\hat{\ell}_{i^{\star},j}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$ (56)
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\ell_{i^{\star},j}\|_{\infty}+2\varepsilon_{\mathsf{stat}}\right)$ (57)
$\leq 2\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\varepsilon_{\mathsf{stat}}.$ (58)
$\leq 2\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\varepsilon_{\mathsf{stat}}.$ (59)
@@
$\lvert J_{M^{\star}}(\pi)-\mathbb{E}_{s\sim d_{0}}[Q_{\hat{i}}(s,\pi)]\rvert$ $\leq 4V_{\max}\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\varepsilon_{\mathsf{stat}}$ (60)
$=24V_{\max}^{3}\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\sqrt{\frac{\log(4dm\delta^{-1})}{n}}.$ (61)
@@ C.3 Proof of Theorem 3
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $=\operatorname{\mathbb{E}}_{d_{0},\pi}\left[Q^{\pi}(s,a)-\widehat{Q}(s,a)\right]$ (62)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\gamma Q^{\pi}(s^{\prime},\pi)-\widehat{Q}(s,a)-\gamma\widehat{Q}(s^{\prime},\pi)\right]$ (63)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)-\widehat{Q}(s,a)+\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)\right]$ (64)
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right]$ (65)
$\leq\frac{1}{1-\gamma}\sqrt{\mathcal{C}^{\pi}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]}$ (66)
where the second line follows from Bellman flow. Now we consider the term under the square root, and let $\widehat{g}_{{\widehat{Q}}}=\operatorname*{arg\,min}_{g\in\mathcal{G}_{\widehat{Q}}}\widehat{\ell}(g,\widehat{Q})$.
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]\leq 2\cdot\underbrace{\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{g}_{{\widehat{Q}}}(s,a)\right)\right]^{2}}_{\mathrm{(T{1})}}+2\cdot\underbrace{\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{\widehat{Q}}}(s,a)-\widehat{Q}(s,a)\right)^{2}\right]}_{\mathrm{(T{2})}}$ (67)
@@
$\widehat{g}_{{Q}}\vcentcolon={}\operatorname*{arg\,min}_{g\in\mathcal{G}_{Q}}\mathbb{E}_{\mathcal{D}}\left[\left(g(s,a)-r-\gamma Q(s^{\prime},\pi)\right)^{2}\right].$ (68)
@@
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q}}(s,a)-\left[\mathcal{T}^{\pi}Q\right](s,a)\right)^{2}\right]\leq\frac{16V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}\vcentcolon={}\varepsilon_{\mathsf{reg}}^{2}.$ (69)
@@
$\max\Bigg{\{}\frac{1}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]~{},$ (70)
$\qquad\quad\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\frac{3}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\Bigg{\}}\leq\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}\vcentcolon={}\varepsilon_{\mathsf{obj}}.$ (71)
@@
[…]
$\leq 2\cdot\varepsilon_{\mathsf{obj}}+2\cdot\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(\widehat{g}_{{\widehat{Q}}}(s,a)-\widehat{Q}(s,a)\right)^{2}\right]$ (74)
$\leq 2\cdot\varepsilon_{\mathsf{obj}}+2\cdot\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-Q^{\pi}(s,a)\right)^{2}\right]$ (75)
$\leq 4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-Q^{\pi}(s,a)\right)^{2}\right]$ (76)
$=4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-[\mathcal{T}^{\pi}Q^{\pi}](s,a)\right)^{2}\right]$ (77)
$\leq 4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\varepsilon_{\mathsf{reg}}^{2}$ (78)
@@
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]$ $\leq 8\left(\varepsilon_{\mathsf{obj}}+\varepsilon_{\mathsf{reg}}^{2}\right),$ (79)

### v3

@@ B.2 Sign-flip Average Bellman Error
$~\mathbb{E}_{\mu}[|Q_{i}-\mathcal{T}_{M^{\star}}^{\pi}Q_{i}|]$
$=$ $~\mathbb{E}_{\mu}[\textrm{sgn}(Q_{i}(s,a)-(\mathcal{T}^{\pi}Q_{i})(s,a))(Q_{i}-\mathcal{T}^{\pi}Q_{i})]$
$=$ $~\mathbb{E}_{\mu}[\textrm{sgn}(Q_{i}-\mathcal{T}^{\pi}Q_{i})(Q_{i}(s,a)-r-\gamma Q_{i}(s^{\prime},\pi))]$
@@
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[Q_{\widehat{i}}(s,\pi)\right]\leq 4\cdot\mathcal{C}_{\infty}^{\pi}\cdot V_{\max}\sqrt{\frac{\log\left(2m/\delta\right)}{n}}.$
@@ C.1 Proof of Theorem 1
$\ell(\theta)\vcentcolon=A\theta-b\in\mathbb{R}^{d},$
$\widehat{\ell}(\theta)\vcentcolon=\widehat{A}\theta-\widehat{b}\in\mathbb{R}^{d}$
@@
$\widehat{\theta}\vcentcolon={}\operatorname*{arg\,min}_{\theta\in\Theta}{}\|\widehat{\ell}(\theta)\|_{\infty}.$
@@
$\|Q^{\pi}(\cdot)-\phi^{\top}(\cdot)\widehat{\theta}\|_{\infty}$ $=\|\phi^{\top}(\cdot)(\widehat{\theta}-\theta^{\star})\|_{\infty}$
$=\|\phi^{\top}(\cdot)A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}(A\widehat{\theta}-b)\|_{\infty}$
[…]
$\leq\left(\max_{s,a}{}\left\|\phi^{\top}(s,a)\right\|_{2}\right)\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\|\ell(\widehat{\theta})\|_{2}$
[…]
@@
[…]
$\leq\|\widehat{\ell}(\theta^{\star})\|_{\infty}+\varepsilon_{\mathsf{stat}}$
$\leq\|\ell(\theta^{\star})\|_{\infty}+2\cdot\varepsilon_{\mathsf{stat}}$
$=2\cdot\varepsilon_{\mathsf{stat}},$
@@
$\|Q^{\pi}-\phi^{\top}\widehat{\theta}\|_{\infty}$ $\leq 2\sqrt{d}B_{\phi}\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\varepsilon_{\mathsf{stat}}$
$=6\sqrt{d}\cdot\left\|A\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$}{}\right\|_{2}\cdot\max\{R_{\max},B_{\phi}\}^{2}\cdot\sqrt{\frac{\log(2d\lvert\Theta\rvert\delta^{-1})}{n}},$
@@
$B^{\pi}(s,a,s^{\prime};{}\theta)\vcentcolon={}\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)$
@@
$\phi_{j}(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)$
$\leq$ $~\|\phi(s,a)\|_{\infty}\left(\|\phi(s,a)\|_{2}\|\theta\|_{2}+\gamma\|\phi(s^{\prime},\pi)\|_{2}\|\theta\|_{2}+R_{\max}\right)$
$\leq$ $~\max_{s,a}\|\phi(s,a)\|_{2}\left(\max_{s,a}\|\phi(s,a)\|_{2}\|\theta\|_{2}+\gamma\max_{s,a}\|\phi(s,a)\|_{2}\|\theta\|_{2}+R_{\max}\right)$
$\leq$ $~(1+\gamma)B_{\phi}^{2}+R_{\max}B_{\phi}$
$\leq$ $~3\max\{B_{\phi},R_{\max}\}^{2}.$
@@
$\left\lvert\operatorname{\mathbb{E}}_{\mu}\left[\phi^{j}(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi^{j}(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]\right\rvert\leq 3\max\{B_{\phi},R_{\max}\}^{2}\sqrt{\frac{2\log(d\lvert\Theta\rvert\delta^{-1})}{n}}=\varepsilon_{\mathsf{stat}},$
@@
$\left\|\ell(\theta)-\widehat{\ell}(\theta)\right\|_{\infty}$ $=\left\|\operatorname{\mathbb{E}}_{\mu}\left[\phi(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi(s,a)\left(\phi^{\top}(s,a)\theta-\gamma\phi^{\top}(s^{\prime},\pi)\theta-r(s,a)\right)\right]\right\|_{\infty}$
$=\left\|\operatorname{\mathbb{E}}_{\mu}\left[\phi(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]-\widehat{\operatorname{\mathbb{E}}}_{\mu}\left[\phi(s,a)B^{\pi}(s,a,s^{\prime};\theta)\right]\right\|_{\infty}$
$\leq\left\|\mathbf{1}\cdot\varepsilon_{\mathsf{stat}}\right\|_{\infty}$
$\leq\varepsilon_{\mathsf{stat}}.$
@@ C.2 Proof of Theorem 2
[…]
@@
[…]
$=\|\phi^{\top}_{\hat{i},i^{\star}}A^{-1}_{\hat{i},i^{\star}}A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{\infty}$
$=\max_{s,a}\lvert\phi^{\top}_{\hat{i},i^{\star}}(s,a)A^{-1}_{\hat{i},i^{\star}}A_{\hat{i},i^{\star}}(e_{2}-e_{1})\rvert$
$\leq\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{2}$
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}(e_{2}-e_{1})\|_{\infty}$
$=\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|A_{\hat{i},i^{\star}}e_{1}-b_{\hat{i},i^{\star}}\|_{\infty}$
$=\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\|\ell_{\hat{i},i^{\star}}\|_{\infty}$
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\|\hat{\ell}_{\hat{i},i^{\star}}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\hat{\ell}_{\hat{i},j}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\hat{\ell}_{i^{\star},j}\|_{\infty}+\varepsilon_{\mathsf{stat}}\right)$
$\leq\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\left(\max_{j\in[m]\setminus\{i^{\star}\}}\|\ell_{i^{\star},j}\|_{\infty}+2\varepsilon_{\mathsf{stat}}\right)$
$\leq 2\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\|A^{-1}_{\hat{i},i^{\star}}\|_{2}\varepsilon_{\mathsf{stat}}.$
$\leq 2\sqrt{d}\left(\max_{s,a}\|\phi_{\hat{i},i^{\star}}(s,a)\|_{2}\right)\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\varepsilon_{\mathsf{stat}}.$
@@
$\lvert J_{M^{\star}}(\pi)-\mathbb{E}_{s\sim d_{0}}[Q_{\hat{i}}(s,\pi)]\rvert$ $\leq 4V_{\max}\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\varepsilon_{\mathsf{stat}}$
$=24V_{\max}^{3}\max_{i\in[m]\setminus\{i^{\star}\}}\frac{1}{\sigma_{\min}(A_{i,i^{\star}})}\sqrt{\frac{\log(4dm\delta^{-1})}{n}}.$
@@ C.3 Proof of Theorem 3
$J(\pi)-\operatorname{\mathbb{E}}_{d_{0}}\left[\widehat{Q}(s,\pi)\right]$ $=\operatorname{\mathbb{E}}_{d_{0},\pi}\left[Q^{\pi}(s,a)-\widehat{Q}(s,a)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\gamma Q^{\pi}(s^{\prime},\pi)-\widehat{Q}(s,a)-\gamma\widehat{Q}(s^{\prime},\pi)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[Q^{\pi}(s,a)-\left[\mathcal{T}^{\pi}Q^{\pi}\right](s,a)-\widehat{Q}(s,a)+\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)\right]$
$=\frac{1}{1-\gamma}\operatorname{\mathbb{E}}_{d^{\pi}}\left[\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right]$
$\leq\frac{1}{1-\gamma}\sqrt{\mathcal{C}^{\pi}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]}$
[…]
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]\leq 2\cdot\underbrace{\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{g}_{{\widehat{Q}}}(s,a)\right)\right]^{2}}_{\mathrm{(T{1})}}+2\cdot\underbrace{\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{\widehat{Q}}}(s,a)-\widehat{Q}(s,a)\right)^{2}\right]}_{\mathrm{(T{2})}}$
@@
$\widehat{g}_{{Q}}\vcentcolon={}\operatorname*{arg\,min}_{g\in\mathcal{G}_{Q}}\mathbb{E}_{\mathcal{D}}\left[\left(g(s,a)-r-\gamma Q(s^{\prime},\pi)\right)^{2}\right].$
@@
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q}}(s,a)-\left[\mathcal{T}^{\pi}Q\right](s,a)\right)^{2}\right]\leq\frac{16V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}\vcentcolon={}\varepsilon_{\mathsf{reg}}^{2}.$
@@
$\max\Bigg\{\frac{1}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]~,$
$\qquad\quad\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]-\frac{3}{2}\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(g(s,a)-Q(s,a)\right)^{2}\right]\Bigg\}\leq\frac{3V_{\max}^{2}\log\left(\frac{2m}{\delta}\right)}{n}\vcentcolon={}\varepsilon_{\mathsf{obj}}.$
@@
$\mathrm{(T2)}$ $=\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{\widehat{Q}}}(s,a)-\widehat{Q}(s,a)\right)^{2}\right]$
$\leq 2\cdot\varepsilon_{\mathsf{obj}}+2\cdot\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(\widehat{g}_{{\widehat{Q}}}(s,a)-\widehat{Q}(s,a)\right)^{2}\right]$
$\leq 2\cdot\varepsilon_{\mathsf{obj}}+2\cdot\operatorname{\mathbb{E}}_{\mathcal{D}}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-Q^{\pi}(s,a)\right)^{2}\right]$
$\leq 4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-Q^{\pi}(s,a)\right)^{2}\right]$
$=4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\operatorname{\mathbb{E}}_{\mu}\left[\left(\widehat{g}_{{Q^{\pi}}}(s,a)-[\mathcal{T}^{\pi}Q^{\pi}](s,a)\right)^{2}\right]$
$\leq 4\cdot\varepsilon_{\mathsf{obj}}+3\cdot\varepsilon_{\mathsf{reg}}^{2}$
@@
$\operatorname{\mathbb{E}}_{\mu}\left[\left(\left[\mathcal{T}^{\pi}\widehat{Q}\right](s,a)-\widehat{Q}(s,a)\right)^{2}\right]$ $\leq 8\left(\varepsilon_{\mathsf{obj}}+\varepsilon_{\mathsf{reg}}^{2}\right),$